# Efficiently Generating Expressive Quadruped Behaviors via Language-Guided Preference Learning

### v1

@@ I INTRODUCTION
The need for user controlled gaits that can rapidly adapt to a user’s preferences has led to a number of recent works that use a natural language due to its intuitiveness and ability to adapt to diverse environments [5, 6, 7, 8]. However, natural language based approaches are fundamentally low-resolution, and it can be challenging for humans to describe the nuanced differences between robot gaits. For example, most humans are not experts on different quadruped locomotion styles. They may not know how to differentiate pacing and trotting, or what different step frequencies look like. So, it can be challenging to specify how a robot should walk without observing and providing specific feedback for existing gaits. On the other hand, preference learning allows users to compare different model outputs to provide high-resolution feedback. This makes it effective for refining robotic behaviors to align with human expectations where direct specification of complex reward functions is impractical [9, 10, 11, 12]. Unfortunately, though learning from preferences allows for greater specificity it comes at a cost — determining a users intent from comparisons alone often requires hundreds [12] or thousands [11] of queries for the most efficient approaches. Querying a human’s feedback hundreds of times to tune the quadruped’s behavior may be tiresome or annoying for the user, and discourage them from adjusting the behavior as needed.
@@ II Related Work
Preference Learning. Due to the challenges in manually specifying rewards for dynamical systems, several techniques have been developed to learn reward functions from human feedback. These methodologies encompass learning from physical corrections [23], scalar feedback [24], rankings [25, 26], and pairwise comparisons [12, 11, 27, 28]. These works often actively generate the next query to ask from the user in order to learn the reward function in a sample efficient manner. Several methods employ preference-based methods to capture complex objectives but often suffer from sample inefficiency [11, 12]. To try and improve sample efficiency, several works develop active learning strategies, some with an LLM in the loop [29, 30], but often struggle in high-dimensional spaces [27, 31]. Moreover, active preference learning is still slow, requiring many samples and multiple feedback iterations. On the other hand real-time systems, like those for quadruped locomotion, necessitate simple and fast feedback mechanisms. Thus, instead of focusing on how to improve feedback over multiple iterations as in active learning, we instead focus on how to generate near-optimal initial reward candidates. In the next section, we show how we can use LLMs to do so efficiently for a real-time quadruped locomotion system. Here we focus on learning a reward vector that is initialized via parameters generated from an LLM given an initial language instruction. We then further adapt such reward functions to align with the non-expert user’s preferences via standard preference learning techniques.

### v2

@@ I INTRODUCTION
The need for user controlled gaits that can rapidly adapt to a user’s preferences has led to a number of recent works that use a natural language due to its intuitiveness and ability to adapt to diverse environments [5, 6, 7, 8]. However, natural language based approaches are fundamentally low-resolution, and it can be challenging for humans to describe the nuanced differences between robot gaits. For example, most humans are not experts on different quadruped locomotion styles. They may not know how to differentiate pacing and trotting, or what different step frequencies look like. So, it can be challenging to specify how a robot should walk without observing and providing specific feedback for existing gaits. On the other hand, preference learning allows users to compare different model outputs to provide high-resolution feedback. This makes it effective for refining robotic behaviors to align with human expectations where direct specification of complex reward functions is impractical [9, 10, 11, 12]. Unfortunately, though learning from preferences allows for greater specificity it comes at a cost — determining a users intent from comparisons alone often requires hundreds [12] or thousands [11] of queries for the most efficient approaches. Querying a human’s feedback hundreds of times to tune the quadruped’s behavior may be tiresome for the user, and discourage them from adjusting the behavior as needed.
@@ II Related Work
Preference Learning. Due to the challenges in manually specifying rewards for dynamical systems, several techniques have been developed to learn reward functions from human feedback. These methodologies encompass learning from physical corrections [23], scalar feedback [24], rankings [25, 26], and pairwise comparisons [12, 11, 27, 28]. These works often actively generate the next query to ask from the user in order to efficiently learn the reward function. Several methods employ preference-based methods to capture complex objectives but often suffer from sample inefficiency [11, 12]. To improve sample efficiency, several works develop active learning strategies, some with an LLM in the loop [29, 30], but often struggle in high-dimensional spaces [27, 31]. Moreover, active preference learning is still slow, requiring many samples and multiple feedback iterations. On the other hand real-time systems, like those for quadruped locomotion, necessitate simple and fast feedback mechanisms. Thus, instead of focusing on how to improve feedback over multiple iterations as in active learning, we instead focus on how to generate near-optimal initial reward candidates. Concurrent to our work [32] also use an LLM provided with in-context information to sample candidate reward functions, which are then evaluated by human users before being used as future in-context examples. In contrast, our work focuses on using preferences to elicit real-time behavior from a fixed multi-task policy.
In the next section, we show how we can use LLMs to do so efficiently for a real-time quadruped locomotion system. Here we focus on learning a reward vector that is initialized via parameters generated from an LLM given an initial language instruction. We then further adapt such reward functions to align with the non-expert user’s preferences via standard preference learning techniques.